A girl with a mass of 65.5 kg is wearing high heeled shoes and standing balanced on one foot.The heel is circular with a diameter of 0.8 cm .What is the pressure exerted by the heel on the horizontal floor ? g=9.8m/s^2


1
Expert's answer
2021-10-01T18:13:04-0400
"p=4\frac{mg}{\pi d^2}\\p=4\frac{65.5\cdot9.8}{\pi (0.008)^2}=1.28\cdot10^7\ Pa"